OpenStudy (agl202):

Quick Help: For the functions f(x) = 2x + 3 and g(x) = 6x + 2, which composition produces the greatest output? Neither composition produces an output. Both compositions produce the same output. f(g(x)) produces the greatest output. g(f(x)) produces the greatest output.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle f(x)=2x+3 }\) \(\large\color{red}{ \displaystyle g(x)=6x+2 }\) \(\large\color{blue}{ \displaystyle f(\color{red}{g(x)})=2\left( \color{red}{6x+2}\right)+3 =? }\)\ \(\large\color{red}{ \displaystyle g(\color{blue}{f(x)})=6\left( \color{blue}{2x+3}\right)+2 =? }\)

OpenStudy (solomonzelman):

I coloered each function in its color, and here belower f(x) and g(x), I am showing how to set up the f(g(x)) and g(f(x))....

OpenStudy (solomonzelman):

this migh be confusing, say so if it is.

OpenStudy (agl202):

What do I do next?

OpenStudy (solomonzelman):

evaluate each composition

OpenStudy (agl202):

how do i do that?

OpenStudy (solomonzelman):

expand the parenthesis in: \(\large\color{blue}{ \displaystyle f(\color{red}{g(x)})=2\left( \color{red}{6x+2}\right)+3 }\)

OpenStudy (solomonzelman):

then simplify as much as you can (by adding like terms)

OpenStudy (agl202):

how do I find x?

OpenStudy (solomonzelman):

you don't need to

OpenStudy (solomonzelman):

just simplify the f(g(x)) and g(f(x)) as much as you can

OpenStudy (solomonzelman):

Again, posting them so that you won't need to scrol way back \(\large\color{blue}{ \displaystyle f(\color{red}{g(x)})=2\left( \color{red}{6x+2}\right)+3 =? }\)\ \(\large\color{red}{ \displaystyle g(\color{blue}{f(x)})=6\left( \color{blue}{2x+3}\right)+2 =? }\)

OpenStudy (agl202):

f(g(x))= 19 g(f(x))= 32 right?

OpenStudy (agl202):

So, D. is the correct choice?

OpenStudy (solomonzelman):

f(g(x)) = 2(6x+2)+3 = 12x+4+3 = 12x + 7 -------------------------------------- g(f(x)) = 6(2x+3)+2=12x+18+2=12x+20 did you plug in 1 for x after that?

OpenStudy (solomonzelman):

But you didn't have that x=1 for that, did you? (Not that it matters..... g(f(x)) is shifted 13 units up from f(g(x)), and for any x it is thus g(f(x)) is greater by 13 units.)

OpenStudy (agl202):

no I didn't.

OpenStudy (solomonzelman):

how did you get 19 and 32 then?

OpenStudy (agl202):

I dunno, how did u get the 4 this equation? --> f(g(x)) = 2(6x+2)+3 = 12x+4+3 = 12x + 7

OpenStudy (solomonzelman):

I expanded, because: